Nam Kading (04.01.2012 to 06.30.2012)

You are here

PDF versionPDF version
Year: 
2012
Quarter: 
April - June
Communications activities: 

 Nothing to report

Unusual Events at the TEAM Site: 

 Nothing to report

New Species: 

 No new species

Protocol Activities: 

Terrestrial Vertebrate - All CT images identified and uploaded to TEAM site.

Vegetation Monitoring - All data uploaded to TEAM site

Climate Monitoring - All data uploaded to TEAM site

Protocol Problems: 

 Climate Monitoring -

1. One Temperature / RH sensor was recieved after calibration and replaced with one needing to be calibrated as per the protocol instructions. All details of the operation were recorded on the TEAM site.

2. We were informed that climatic data from the month of June 2011 was missing, and various steps were taken to extract this missing data, but without success. On 18 June 2012, we downloaded all stored data directly from the logger and submitted to TEAM fo see if they were able to extract the missing data from these files.

3. The Precipitation gauge was found to be under-reading the rainfall in comparison to the manual rain gauge. Investigation revealed that the brass siphon valve was clogged and after cleaning, was found to be operating normally. This maintanance item was added to the maintenance schedule in more detail and the staff working at the climate station were given onsite training on how to carry out this task. This took place on the 18 June 2012, so precipitation data for a month prior to this may be inacurate.   

Schedule Problems: 

 Nothing to report

Logistical Problems: 

 Nothing to report

Suggestions: 

 Nothing to report